Express the following polynomial as a linear combination of the polynomials.


P=t2+ 4t-3

P1=t2+2t+5

P2=2t2-3t

P3=t+3


1
Expert's answer
2021-06-29T12:34:08-0400

Let us express the following polynomial as a linear combination of the polynomials.


"P=aP_1+bP_2+cP_3"


It follows that


"t^2+ 4t-3= a(t^2+2t+5)+b(2t^2-3t)+c(t+3)",


and hence


"t^2+ 4t-3= (a+2b)t^2+(2a-3b+c)t+(5a+3c)"


We get the following system:


"\begin{cases} a+2b=1\\2a-3b+c=4\\5a+3c=-3\end{cases}"


"\begin{cases} a=1-2b\\2(1-2b)-3b+c=4\\5(1-2b)+3c=-3\end{cases}"


"\begin{cases} a=1-2b\\-7b+c=2\\-10b+3c=-8\end{cases}"


"\begin{cases} a=1-2b\\c=2+7b\\-10b+3(2+7b)=-8\end{cases}"


"\begin{cases} a=1-2b\\c=2+7b\\11b=-14\end{cases}"


"\begin{cases} a=\frac{39}{11}\\c=-\frac{76}{11}\\b=-\frac{14}{11}\end{cases}"


We conclude that


"P=\frac{39}{11}P_1-\frac{14}{11}P_2-\frac{76}{11}P_3"